OpenStudy (anonymous):

|x+3|=5 what is the answer to this absolute value question?

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

absolute value just means there r no negatives. so its basically just like solving x+3=5 -3 -3 x=2

OpenStudy (anonymous):

x=2 and -2 actually c:

OpenStudy (anonymous):

x is either 2 or -8, because you want the expression inside the absolute value to be 5 OR -5.

OpenStudy (anonymous):

can you please show the exact steps?

OpenStudy (anonymous):

x+3=5 - 3 -3 --- x=2

OpenStudy (anonymous):

\[|x+3|=5\] -3 -3 x = 2

OpenStudy (anonymous):

Guys, all absolute value equations have 2 answers... -8 + 3 = -5, and \[\left| -5 \right|\] = 5, thus -8 works here, as well as 2.

hero (hero):

What the...@bieberfan is still around?